Please help me solve the following problem:
The instructions read: List the given point and slope of y + 1 = 4/5 (x + 4)
I keep getting fractions that do not work when I place them back into the equation.
Thank you

Answer by stanbon(75887)

Please help me solve the following problem:
The instructions read: List the given point
and slope of y + 1 = 4/5 (x + 4)
---------------
Solve for "y":
y+1 = (4/5)(x+4)
y+1 = (4x/5)+(16/5)
---
y = (4/5)x +11/5
----
slope = 4/5
y-intercept = 11/5
========================
Cheers,
Stan H.